3(x+2.86)=2x Ive done this question a million times cant figure out where I messed up
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

3(x+2.86)=2x

3x+3(2.86)=2x

3x+8.58=2x

8.58=2x-3x

8.58=-x

-x = 8.58

x = -8.58

So the solution is x = -8.58

Answer by stanbon(75887)

3(x+2.86)=2x
------
x + 2.86 = (2/3)x
-----
(1/3)x = -2.86
---
x = 3(-2.86)
----
x = -8.58
=================
Cheers,
Stan H.